OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 15x^4+ 50x^3– 40x^2.

OpenStudy (anonymous):

=5x^2(3x^2+10x-8)=5x^2(3x^2+12x-2x-8)=5x^2[3x(x+4)-2(x+4)]=5x^2(3x-2)(x+4)